# Non-perturbative approach for scalar particle production in Higgs-$R^{2}$ inflation

### v1

@@ V Description of particle production
This behavior usually happens at early times $t_{i}\to 0$ and at sufficiently late times $t_{f}\to+\infty$, but for intermediate times $t_{i}<t_{*}<t_{f}$ during the expansion of the universe the evolution of the modes is no longer adiabatic $|A_{k}|\gg 1$. During inflation, the positive-frequency solution corresponds to the Bunch-Davies vacuum (45), as described in the previous section. This evolution remains valid as long as the parameter $|A_{k}|\ll 1$ at all times. While this condition is met, no particle production occurs, as adiabatic evolution is ensured, and there is no mixing of negative- and positive-frequency modes. From figure 3, we can see that the evolution of $\omega_{k}(t)$ is smooth until the end of inflation. At that point, the condition $|A_{k}|\ll 1$ is strongly violated, leading to an enhancement in particle production and gravitational particle production is enhanced at the time $t_{*}$, which corresponds to the end of inflation. We show the non-adiabaticity evolution in the figure 4 parametrized by the adiabatic parameter $|A_{k}(t)|$. The mixture of positive and negative frequency modes at time $t_{\mathrm{end}}$ can be represented as the sum of the solutions $\exp(\pm i\int^{t}\omega_{k}(t^{\prime})\mathrm{d}t^{\prime})$ using the WKB approximation [14, 5]
@@
where $\alpha_{k}(t)$ and $\beta_{k}(t)$ are the time-dependent Bogolyubov coefficients. Comparing with the Bunch-Davies initial condition (47), one see thats $\alpha_{k}(t_{0})=1$ and $\beta_{k}(t_{0})=0$. The comoving particle number density $n_{\chi}(t)$ is related to the coefficient $\beta_{k}$ of the Bogolyubov transformations which relate the vacuum states at the asymptotic limits $t\to 0$ and $t\to\infty$, and coincides with $\beta_{k}(t)$ at late times. We can compute $\beta_{k}(t)$ by solving (40) and evaluating
@@
Figure 5: Left: Comoving number density of particles created for $m_{\chi}=10^{-1}M$ and $\xi_{\chi}=1/6$. The graph shows the evolution of the occupation number $n_{k}/k_{\mathrm{end}}^{3}$ as a function of time $t$, where the mode $k$ was chosen to be 5 e-folds before leaving the horizon during inflation and $k_{\mathrm{end}}=a_{\mathrm{end}}H_{\mathrm{end}}$ is the mode at the end of inflation. Right: Spectrum of the gravitationally produced scalar particles for the same numerical example. Modes with $k=aH$ are on the horizon scale at the end of inflation. The mass and the momenta $k$ both are given in units of $M$.
@@
where $n_{k}(t)$ is the comoving number density of particles with comoving wavenumber $k$.
V.1 Numerical results
@@
What follows is to find the spectrum of $\chi$ particles and its evolution in the cosmic time $t$. Hence we numerically solve (40), constructed the Bogolyubov coeffcient $|\beta_{k}(t)|$ and compute the spectrum $n_{k}$ using (50). In the previous section we described the procedure to numerically evaluate the set of equations to ensure that each mode $k$ is within the horizon at the beginning of the evolution, which allows us to use the Bunch-Davies vacuum (45) as a initial condition of the mode equation for $X_{K}(t)$. The initial conditions for the background are set at time $t_{\mathrm{ini}}$, corresponding to the moment when the modes are within the horizon. For example, for $N_{*}=60$ at the time of crossing, the initial time corresponds to $t_{\mathrm{ini}}\approx 28.712$ in units of $[10^{-5}M_{p}^{-1}]$, and the corresponding mode $k=aH$ will be the value of $aH$ at time $t_{\mathrm{ini}}$. Then, the Bogolyubov coefficient $|\beta_{k}(t)|^{2}$ is calculated using (49) and the comoving number density of gravitational particle production is determined by $n_{k}(t)$.
V.1.1 Conformal coupling $\xi_{\chi}=1/6$
@@
First of all, we focus on the conformal coupling $\xi_{\chi}=1/6$ and consider light masses $m_{\chi}\ll M/2$. As first example, we show the the comoving number density of particles $n_{k}(t)$ as a function of time $t$ and the corresponding spectrum for the case of $m_{\chi}=10^{-1}M$. In the left panel of figure 5, we observe that the number of particles created saturates at the end of inflation, stabilizing and becoming constant at late times $t\to\infty$. This growth in $n_{k}$ is due to the non-adiabaticity of $\omega_{k}$ at the end of inflation, as can be seen in the figure 4. The oscillations of $n_{k}(t)$ after inflation are due to resonances at certain moments between the oscillations of the modes $X_{k}$ and the frequency itself. In the right panel, we observe the corresponding spectrum of $n_{k}(t)$. Several remarkable features can be observed. For IR modes $(k\ll aH)$, we see that the spectrum is blue-tilted, reaching a peak at some intermediate value $k_{\mathrm{peak}}$, and then is suppressed for UV modes $(k\gg aH)$ due to inflation, that is, those that remain within the horizon during inflation. Modes with $k\ll k_{\mathrm{end}}$ leaves the horizon during inflation and re-enter later during the reheating epoch after inflation. Therefore, the dynamics of the $X_{k}(t)$ modes are dictated by the effective mass term (43) with a quasi-De Sitter spacetime as background with $H\simeq\mathrm{constant}$. It is important to note that the UV modes cross the horizon just at the end of inflation, leading to the production of these modes at that instant and their subsequent evolution during the preheating epoch after inflation, as seen in the left panel of figure 5. During this epoch, the production occurs mainly due to the oscillatory nature of the scalaron and the Higgs field, as shown in figure 2, through the term $\dot{H}$ and the Ricci scalar $R$ via the non-minimal coupling $\xi_{\chi}=1/6$. The suppression of UV modes occurs because the $\chi$ scalars become supermassive at the end of inflation due to the exponentially scalaron-dependence mass, causing the effective mass to become substantial. As a result, gravitational particle production becomes less efficient, leading to a significant suppression in the particle density number $n_{k}(t)$.
Figure 6 presents the spectrum of the comoving number of particles produced across the range of light masses considered in this study.
Figure 6: Comoving spectrum of gravitational scalar particle production $n_{k}$ for a range of light masses of the field $\chi$ as a function of rescaled horizon modes $k/k_{\mathrm{end}}$. Masses are given in units of scalaron mass $M$. One can seen that the UV tail is suppressed for light masses $m_{\chi}<M/2$.
The numerical results exhibit a blue-tilted behavior in the infrared (IR) modes, with an exponential suppression in the ultraviolet (UV) tail for light masses $m_{\chi}<M/2$. The mass of the field plays a key role in determining the amplitude of $n_{k}$ at the peak momentum $k_{\mathrm{peak}}$, which systematically decreases as $m_{\chi}$ increases [44]. It is important to emphasize that, in this mass range, particle production following inflation is predominantly driven by the decay of the scalaron into pairs of light $\chi$ particles, as described by expression (33). Furthermore, figure 6 indicates that non-perturbative particle production within this mass range remains minimal, suggesting that such production could lead to an overproduction [8, 50]. For the specific cases of $m_{\chi}=0.2M$ and $m_{\chi}=0.3M$, the production is significantly more moderate. The spectra also exhibit a tendency to flatten in the UV region, implying that gravitational particle production becomes increasingly inefficient for larger masses.
V.1.2 Non-minimal coupling $\xi_{\chi}=0$
@@
Figure 7: Comoving spectrum of gravitational scalar particle production $n_{k}$ for $\xi_{\chi}=0$ as a function of rescaled horizon modes $k/k_{\mathrm{end}}$. Masses are given in units of scalaron mass $M$. One can seen that the spectrum is nearly scale-invariant with a blue-tilted behavior for larger masses.
Figure 8: Comoving number density of particles $n_{k}/k_{\mathrm{end}}^{3}$ for $m_{\chi}=10^{-1}M$ and $\xi_{\chi}=0$ as a function of cosmic time $t$.
@@
For the case of $\xi_{\chi}=0$, we have a different situation. During inflation, the effective mass (43) can be negative for light masses $m_{\chi}\ll H_{\mathrm{end}}$
$m_{\mathrm{eff}}^{2}(t)\approx e^{-\alpha\phi}m_{\chi}^{2}-\dfrac{9}{4}H^{2}-\dfrac{\alpha}{2}V_{\phi}<0\,,$
and once the modes cross the horizon, we obtain $\omega_{k}^{2}(t)<0$, leading to tachyonic instabilities as pointed out in the discussion following equation (43). Therefore, particle production can be due to resonances from oscillations or due to the instabilities caused by $\omega_{k}^{2}(t)<0$. In figure 8, we can observe the behavior of $n_{k}(t)$ normalized to its value at the end of inflation as a function of cosmic time. We can see that for the same mass value $m_{\chi}=10^{-1}M$, particle production is significantly larger than in the case of conformal coupling (see figure 5).
The tachyonic growth of the modes $X_{k}(t)$ outside the horizon is the main driver for the enhanced production of minimally coupled particles compared to the $\xi_{\chi}=1/6$ case. We observe that after inflation, $n_{k}(t)$ exhibits an anharmonic oscillatory behavior with maximum production peaks until it reaches a constant value as $t\to\infty$. However, for masses $m_{\chi}\gg H_{\mathrm{end}}$, the dominant terms in the effective mass during inflation are $e^{-\alpha\phi}m_{\chi}^{2}$ and $(k/a)^{2}$, which implies a positive frequency $\omega_{k}^{2}(t)>0$. In figure 7, we show the spectrum of scalar particles created in the case of minimal coupling and masses $m_{\chi}\gtrsim M/2$. As can be seen, the spectrum generated in the IR region is qualitatively different as the mass increases. We observe an almost scale-invariant spectrum that is slightly blue-tilted for masses $m_{\chi}>M/2$.
VI Discussion and remarks
@@
In this study, we investigated the gravitational and non-perturbative production of free scalar particles, $\chi$, in Higgs-$R^{2}$ inflation. We began with a free spectator field, $\chi$, interacting solely through the gravitational term $\xi_{\chi}R$ in the Jordan frame. By applying a conformal transformation to transition into the Einstein frame, direct couplings between the $\chi$ field and the scalaron $\phi$ emerged. Within the context of inflationary dynamics in the Einstein frame, we derived the equation of motion for the $\chi$ field, taking into account the non-minimal coupling with R. We numerically solved this system of equations, imposing Bunch-Davies conditions on the $\chi$ field 5 e-folds before the modes crossed the horizon, to determine the mode functions of the $\chi$ field. From these solutions, we extracted the Bogolyubov coefficient $\beta_{k}(t)$ to calculate the comoving particle density and visualize their spectrum. We explored two distinct scenarios for the non-minimal coupling $\xi_{\chi}$, considering different mass ranges. For $\xi_{\chi}=1/6$ and light masses, we observed efficient particle production at the end of inflation due to the term $e^{-\alpha\phi}m_{\chi}^{2}$, which becomes dominant at that stage. The resulting power spectrum exhibited a blue-tilted behavior, indicating that particle production was dominated by modes near the Hubble radius $(aH)^{-1}$ at the end of inflation, while shorter wavelength modes were suppressed.
@@
In the case of $\xi_{\chi}=1/6$ and for masses close to $m_{\chi}\sim M/2$, the spectrum tends to flatten out without exhibiting a peak in the UV region. This phenomenon is possibly due to the mass term that depends exponentially on the scalaron $\phi$, causing the effective mass to become superheavy at the end of inflation. In contrast, for $\xi_{\chi}=0$, the spectrum remained nearly scale-invariant for masses $m_{\chi}\gtrsim M/2$, with a blue tilt emerging for larger masses. For light masses and $\xi_{\chi}=0$, we find that the comoving number of particles created is significantly higher compared to the case of conformally coupled particles, which is due to the tachyonic growth of the modes $X_{k}(t)$. We have not attempted a detailed exploration of the parameter space of the Higgs-$R^{2}$ inflation model or the parameters of the scalar field $\chi$. Instead, we have taken the Higgs-$R^{2}$ inflation model, whose observables are consistent with Planck, and considered the possibility of non-perturbative particle production through the direct couplings that arise between the $\chi$ field and the inflationary sector when changing the conformal frame. We found that inflation driven by the Higgs field and the scalaron can produce scalar particles in the cases $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$ and for a specific range of masses. This field could potentially contribute to the abundance of dark matter, for which its current abundance would need to be calculated.
There remains much to be done for future work, such as considering the production of spin-$1/2$ and spin-$1$ dark matter, the production of gauge bosons $W^{\pm}_{\mu}$, $Z_{\mu}$, as well as the implications on primordial non-Gaussianities in the CMB induced by particle production. It would also be interesting to study the isocurvature constraints for light scalar field masses in this model. We leave these intriguing topics for future studies.
VII Acknowledgements

### v2

@@ 5 Description of particle production
This behavior usually happens at early times $t_{i}\to 0$ and at sufficiently late times $t_{f}\to+\infty$, but for intermediate times $t_{i}<t_{*}<t_{f}$ during the early universe, the evolution of the modes is no longer adiabatic $|A_{k}|\gg 1$. During inflation, the positive-frequency solution corresponds to the Bunch-Davies vacuum (43), as described in the previous section. This evolution remains valid as long as the parameter $|A_{k}|\ll 1$ at all times. While this condition is met, no particle production occurs, as adiabatic evolution is ensured, and there is no mixing of negative- and positive-frequency modes. From figure 3, we can see that the evolution of $\omega_{k}(t)$ is smooth until the end of inflation. At that point, the condition $|A_{k}|\ll 1$ is strongly violated, leading to an enhancement in particle production, and gravitational particle production is enhanced at the time $t_{*}$, corresponding to the end of inflation. We show the non-adiabaticity evolution in the figure 4 parametrized by the adiabatic parameter $|A_{k}(t)|$. The mixture of positive and negative frequency modes at time $t_{\mathrm{end}}$ can be represented as the sum of the solutions $\exp(\pm i\int^{t}\omega_{k}(t^{\prime})\mathrm{d}t^{\prime})$ using the WKB approximation [14, 5]
@@
where $\alpha_{k}(t)$ and $\beta_{k}(t)$ are the time-dependent Bogolyubov coefficients. Comparing with the Bunch-Davies initial condition (45), one see thats $\alpha_{k}(t_{0})=1$ and $\beta_{k}(t_{0})=0$.
Figure 4: An example of non-adiabatic mode evolution leading the gravitational particle production of $\chi$ particles described by the adiabaticity parameter $|A_{k}(t)|$ for $m_{\chi}=10^{-2}H_{\mathrm{end}}$ and $\xi_{\chi}=1/6$. At the end of inflation, one sees an abrupt increase on $|A_{k}(t)|$ and non-adiabaticity is induced, leading to gravitational particle production $n_{k}(t)\neq 0$. At late times, when the stabilization occurs, we recover the adiabaticity evolution.
In other words, gravitational particle production occurs at very early times $t_{\mathrm{in}}\to 0$ and at late times $t_{f}\to\infty$ if the initial and final vacuum states are not the same. We define the initial vacuum state with the corresponding creation-annihilation operators $\hat{a}^{\dagger}_{\mathbf{k}}$ and $\hat{a}_{\mathbf{k}}$ by
$\hat{a}_{\mathbf{k}}\left|0\right>_{\mathrm{early}}=0\,.$ (47)
The final state will have a different set of creationist-annihilation operators, and thus a different vacuum state. $\hat{a}_{\mathbf{k}}$ it is related to the late-time operators $\hat{b}^{\dagger}_{\mathbf{k}}$ and $\hat{b}_{\mathbf{k}}$ by a Bogolyubov tranformation
$\hat{a}_{\mathbf{k}}=\alpha^{*}_{\mathrm{k}}\hat{b}_{\mathbf{k}}-\beta^{*}_{\mathrm{k}}\hat{b}^{\dagger}_{\mathrm{k}}\,,$ (48)
where $\alpha_{\mathbf{k}}$ and $\beta_{\mathbf{k}}$ are the time-independent Bogolyubov coefficients. The comoving particle number density $n_{\chi}(t)$ is related to the coefficient $\beta_{k}$ of the Bogolyubov transformations which relate the vacuum states at the asymptotic limits $t\to 0$ and $t\to\infty$ and coincides with $\beta_{k}(t)$ at late times. We can compute $\beta_{k}(t)$ by solving (38) and evaluating
@@
where $n_{k}(t)$ is the comoving number density of particles with comoving wavenumber $k$. Notice that the expression (49) is only valid when $\omega_{k}^{2}(t)>0$, so the comoving number density of particles created is valid at times when the frequency (39) is positive.
5.1 Numerical results
@@
What follows is to find the spectrum of $\chi$ particles and their evolution in the cosmic time $t$. Hence we numerically solve (38), constructed the Bogolyubov coefficient $|\beta_{k}(t)|$, and compute the spectrum $n_{k}$ using (50). In the previous section, we described the procedure to numerically evaluate the set of equations to ensure that each mode $k$ is within the horizon at the beginning of the evolution, which allows us to use the Bunch-Davies vacuum (43) as an initial condition of the mode equation for $X_{K}(t)$. The initial conditions for the background are set at time $t_{\mathrm{ini}}$, corresponding to the moment when the modes are within the horizon. For example, for $N_{*}=60$ at the time of crossing, the initial time corresponds to $t_{\mathrm{ini}}\approx 28.712$ in units of $[10^{-5}M_{p}^{-1}]$, and the corresponding mode $k=aH$ will be the value of $aH$ at time $t_{\mathrm{ini}}$. Then, the Bogolyubov coefficient $|\beta_{k}(t)|^{2}$ is calculated using (49), and the comoving number density of gravitational particle production is determined by $n_{k}(t)$.
5.1.1 Conformal coupling $\xi_{\chi}=1/6$
We first present the results for conformally coupled particles ($\xi_{\chi}=1/6$). As an example, we present the time evolution of the comoving number of particles created $n_{k}(t)$ as a function of time for two different cases: light particles $m_{\chi}<H_{\mathrm{end}}$ and heavy particles $m_{\chi}\gtrsim H_{\mathrm{end}}$. We should note that the expression for $\beta_{k}(t)$ is only valid when $\omega_{k}^{2}(t)>0$ , we point out that $n_{k}(t)$ only makes physical sense at late times $t\to\infty$ [5]. The frequency (39) for scalar conformally coupled turns out to be
$\omega_{k}^{2}(t)=\left(\dfrac{k}{a}\right)^{2}+e^{-\alpha\phi}m_{\chi}^{2}+\dfrac{R}{6}-\dfrac{9}{4}H^{2}-\dfrac{3}{2}\dot{H}$ (51)
which can be negative for $k\ll aH$. For this reason, we take into account the sub-horizon evolution of the modes $X_{k}(t)$ and we take $k$ deep inside the horizon at the beginning of the simulation. In the figure, we present two cases: $m_{\chi}=10^{-5}H_{\mathrm{end}}$ and $m_{\chi}=10^{-1}H_{\mathrm{end}}$. In both panels of figure 5, we observe that the number of particles created saturates at the end of inflation, stabilizing and becoming constant at late times $t\to\infty$. This growth in $n_{k}$ is due to the non-adiabaticity of $\omega_{k}$ at the end of inflation, as can be seen in figure 4.
Figure 5: Comoving number density of particles, plotted on a logarithmic scale, as a function of cosmic time $t$ for two different values of the spectator mass $m_{\chi}$, with the same mode $k$, corresponding to $N_{k}=60$. The graph depicts the evolution of the occupation number $n_{k}/k_{\mathrm{end}}^{3}$, where $k_{\mathrm{end}}=a_{\mathrm{end}}H_{\mathrm{end}}$ denotes the mode at the end of inflation
The oscillations of $n_{k}(t)$ after inflation arise from resonances occurring at specific moments between the oscillations of the modes $X_{k}$ and the frequency itself. During this epoch, particle production is driven primarily by the oscillatory behavior of the scalaron and the Higgs field, as illustrated in figure 2, through the term $\dot{H}$ and the Ricci scalar $R$ via the non-minimal coupling $\xi_{\chi}=1/6$. Additionally, we observe that particle production near the end of inflation is most efficient for light particles. In both cases, the particle number is well-defined, grows from the end of inflation, and stabilizes at late times, as expected.
Figure 6: Numerical $n_{k}/k_{\mathrm{end}}^{3}$ vs $k/k_{\mathrm{end}}$ spectra of conformally-coupled scalars for differente light masses $m_{\chi}$. One can see that the UV tail is suppressed for this range of masses.
Figure 7: Right: Comoving number density of particles as a function fo cosmic time $t$. Left: Numerical $n_{k}/k_{\mathrm{end}}^{3}$ vs $k/k_{\mathrm{end}}$ spectra of conformally-coupled scalars for differente large masses $m_{\chi}\gtrsim H_{\mathrm{end}}$. One can see that the UV tail is suppressed for this range of masses.
@@ 5.1.1 Conformal coupling $\xi_{\chi}=1/6$
We have numerically computed the spectra of the comoving number density $n_{k}(t)$ as a function of $k/k_{\mathrm{end}}$ for a range of light masses varying from $10^{-5}H_{\mathrm{end}}$ to $10^{-1}H_{\mathrm{end}}$. Our findings are summarized in figure 6. From the figure, one can observe several remarkable features. Conformally coupled scalars show a spectra that are strongly peaked at some value $k_{\mathrm{peak}}$. For IR modes $(k\ll k_{\mathrm{end}})$, we see that the spectrum is blue-tilted, reaching a peak at some intermediate value $k_{\mathrm{peak}}>k_{\mathrm{end}}$, and then is suppressed for UV modes $(k\gg k_{\mathrm{end}})$ due to inflation, that is, those that remain within the horizon during inflation. IR modes leave the horizon during inflation and re-enter later during the reheating epoch after inflation. Therefore, the dynamics of the $X_{k}(t)$ modes are dictated by the effective mass term (41) with a quasi-De Sitter spacetime as background with $H\simeq\mathrm{constant}$. It is important to note that the UV modes do not cross the horizon during inflation. In place, these modes are mainly produced during the transition to the matter-dominated era and in general during the reheating era. The suppression of UV modes occurs because the $\chi$ scalars become supermassive at the end of inflation due to the exponentially scalaron-dependence mass, causing the effective mass to become substantial. As a result, gravitational particle production becomes less efficient, leading to a significant suppression in the particle density number $n_{k}(t)$. The mass of the field plays a key role in determining the amplitude of $n_{k}$ at the peak momentum $k_{\mathrm{peak}}$, which systematically decreases as $m_{\chi}$ increases [43].
For large masses $m_{\chi}\gtrsim H_{\mathrm{end}}$, the situation is quite similar. As the mass increases, the particle production tends to decrease due to the increase of the effective mass. This can be seen in the left panel of figure 7 for a mass of $m_{\chi}=H_{\mathrm{end}}$ and for a mode crossing the horizon at $N_{k}=60$. In the rigt panel, we present the comoving spectra of particle production for $m_{\chi}\gtrsim H_{\mathrm{end}}$. We find that the behaviour for UV modes is different than in the case of light masses. The location of the peak $k_{\mathrm{peak}}$ is different as a function of the mass value, once $k_{\mathrm{end}}$. While $k_{\mathrm{peak}}>k_{\mathrm{end}}$ as in the case of light particles, the entry to the horizon, and hence the tail suppression of the UV modes is different. This is largely due to the possible resonances that exist for large masses, as the mass term $e^{-\alpha\phi}$ dominates after inflation when $\phi<M_{p}$, which can excite different modes after inflation.
@@
Figure 8: The comoving spectrum of gravitational scalar particle production $n_{k}$ for $\xi_{\chi}=0$ as a function of rescaled horizon modes $k/k_{\mathrm{end}}$. One can see that the spectrum is nearly scale-invariant for masses $m_{\chi}\lesssim H_{\mathrm{end}}$ and blue-tilted behavior for larger masses
5.1.2 Minimal coupling $\xi_{\chi}=0$
@@
For the case of $\xi_{\chi}=0$, we have a different situation. During inflation, the frequency (39) is given by
$\omega_{k}^{2}(t)=\left(\dfrac{k}{a}\right)^{2}+e^{-\alpha\phi}m_{\chi}^{2}-\dfrac{9}{4}H^{2}+\dfrac{7}{6}\dot{H}-\dfrac{\alpha}{2}V_{\phi}\,.$ (52)
In this case, the derivative potential $V_{\phi}$ and the Hubble parameter $H^{2}$ contribute so that the effective mass $m_{\mathrm{end}}$ is negative. Thus, once the modes cross the horizon and for masses $m_{\chi}\ll H_{\mathrm{end}}$, this may lead to $\omega_{k}^{2}(t)<0$, resulting in tachyonic instabilities, as mentioned in the discussion following equation (41) [41]. These instabilities primarily affect the IR modes of the produced particle spectrum, as the modes outside the horizon during inflation experience tachyonic growth. This growth of the modes $X_{\mathbf{k}}(t)$ outside the horizon is the main driver of the enhanced particle production for minimally coupled scalars compared to conformally coupled scalars. The spectrum of the produced particles is also affected, and its behavior depends on the mass $m_{\chi}$. To obtain the gravitational spectra, we use the same technique to solve the equation numerically: we choose the $k$-mode inside the horizon, approximately 5 e-folds before it crosses the horizon and show the numerical result for $\beta_{k}(t)$ at those times when $\omega_{k}^{2}>0$. In the figure 8, we show the spectrum of created particles for minimally coupled masses. We observe that the spectrum is nearly scale-invariant for IR modes and for masses $m_{\chi}\lesssim H_{\mathrm{end}}$ before reaching the peak at $k_{\mathrm{peak}}$. However, as the mass increases, the spectrum tends to be blue-tilted. In the UV region, all modes decay either as a power law or exponentially.
Figure 9: Comoving number density of particles as a function of cosmic time. A similar behavior can be observed in Figure 6: the number of created particles saturates at the end of inflation and stabilizes, becoming constant at late times
A clear difference exists between the cases $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$. In the frequency $\omega_{k}^{2}(t)$, the Ricci scalar does not appear for $\xi_{\chi}=0$, indicating a significant difference in the behavior of $\omega_{k}$ for different coupling cases. This is because the Ricci scalar exhibits an oscillatory behavior after inflation due to the oscillations of the fields $\phi$ and $h$. In other words, for the case $\xi_{\chi}=1/6$, the Ricci scalar plays an important role in describing the evolution of the modes $X_{k}(t)$, which translates into a different spectrum of created particles. From figure 9 we can see that the particle production at minimal coupling is much larger than the conformally-coupled case. This is mainly due to the behavior of the spectator field $X_{k}(t)$ during inflation. As explained above, the tachyonic instabilities in the case of $\xi_{\chi}=0$ cause the modes $X_{k}(t)$ to grow exponentially. This growth of the super-horizon modes is primarily responsible for the enhanced gravitational particle production for minimally coupled scalars.
6 Discussion and remarks
@@
In this study, we investigated the gravitational and non-perturbative production of free scalar particles, $\chi$, in Higgs-$R^{2}$ inflation. We started with a free spectator field, $\chi$, interacting only through the gravitational term $\xi_{\chi}R$ in the Jordan frame. By applying a conformal transformation to transition into the Einstein frame, direct couplings between the $\chi$ field and the scalaron $\phi$ naturally emerged. Within the context of inflationary dynamics in the Einstein frame, we derived the equation of motion for the $\chi$ field, accounting for its non-minimal coupling to the scalar curvature $R$. We numerically solved this system of equations, imposing Bunch-Davies conditions on the $\chi$ field five e-folds before the modes crossed the horizon, to determine the mode functions of $\chi$. From these solutions, we extracted the Bogolyubov coefficient $\beta_{k}(t)$ to compute the comoving particle density and visualize the resulting spectrum. We explored two distinct scenarios for the non-minimal coupling parameter $\xi_{\chi}$, considering different mass ranges. For $\xi_{\chi}=1/6$, we observed efficient particle production at the end of inflation, driven by the term $e^{-\alpha\phi}m_{\chi}^{2}$, which becomes dominant at this stage. The resulting power spectrum exhibited a blue-tilted behavior, indicating that particle production was dominated by modes near the Hubble radius $(aH)^{-1}$ at the end of inflation, while shorter-wavelength modes were suppressed. As the mass increased in the $\xi_{\chi}=1/6$ case, particle production tended to decrease due to the increase in the effective mass. In contrast, for $\xi_{\chi}=0$, the spectrum remained nearly scale-invariant for masses $m_{\chi}\lesssim H_{\mathrm{end}}$, with a blue tilt emerging for larger masses $m_{\chi}\gtrsim H_{\mathrm{end}}$. For light masses and $\xi_{\chi}=0$, we found that the comoving number of particles created was significantly higher compared to the case of conformally coupled particles, primarily due to the tachyonic growth of the modes $X_{k}(t)$.
@@
We did not attempt a detailed exploration of the parameter space of the Higgs-$R^{2}$ inflation model or the parameters of the scalar field $\chi$. Instead, we focused on the Higgs-$R^{2}$ inflation model, whose observables are consistent with Planck data, and considered the possibility of non-perturbative particle production through the direct couplings that arise between the $\chi$ field and the inflationary sector when changing the conformal frame. We found that inflation driven by the Higgs field and the scalaron can produce scalar particles in the cases $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$ for a specific range of masses. This field could potentially contribute to the abundance of dark matter, though its present-day abundance would need to be calculated. There remain several directions for future research, including the production of spin-$1/2$ and spin-$1$ dark matter, the production of gauge bosons $W^{\pm}_{\mu}$, $Z_{\mu}$, and the implications of particle production on primordial non-Gaussianities in the CMB. It would also be interesting to study the isocurvature constraints for light scalar field masses in this model. We leave these intriguing topics for future investigation.
[…]
Acknowledgements